OpenStudy (anonymous):

how do you find the area of a circle with a circumference of 42

OpenStudy (amistre64):

area = pi(r^2) circumference = 2pi(r) solve the circumference for "r" and use it in the area...

OpenStudy (amistre64):

Do you mean 42pi for circumference? or just 42? 42 = 2pi(r) r = 21/pi A = pi(2/pi)^2 A = pi(2)(2) ------ pi(pi) A = 4/pi

OpenStudy (amistre64):

42pi = 2pi(r) r = 21 A = pi(r^2) A = pi(21^2) A = 441..if I did it right in me head